### v1

# Efficient Time Series Processing for Transformers and State-Space Models through Token Merging

Abstract
Transformer architectures have shown promising results in time series processing. However, despite recent advances in subquadratic attention mechanisms or state-space models, processing very long sequences still imposes significant computational requirements. Token merging, which involves replacing multiple tokens with a single one calculated as their linear combination, has shown to considerably improve the throughput of vision transformer architectures while maintaining accuracy. In this work, we go beyond computer vision and perform the first investigations of token merging in time series analysis on both time series transformers and state-space models. To effectively scale token merging to long sequences, we introduce local merging, a domain-specific token merging algorithm that selectively combines tokens within a local neighborhood, adjusting the computational complexity from linear to quadratic based on the neighborhood size. Our comprehensive empirical evaluation demonstrates that token merging offers substantial computational benefits with minimal impact on accuracy across various models and datasets. On the recently proposed Chronos foundation model, we achieve accelerations up to with only minor accuracy degradations.
1 Introduction
Since their inception in NLP [1], transformers have extended their influence into various domains, including computer vision with Vision Transformers (ViTs) [2], graphs [3, 4], and time series processing [5]. However, the computational complexity of the standard attention mechanism used in transformer architectures scales quadratically with the number of input tokens, resulting in high memory requirements. This scalability issue becomes especially pronounced in time series processing, where sequences frequently comprise thousands of tokens [6]. Consequently, recent foundational models in time series, such as Chronos, exhibit impressive zero-shot generalization capabilities but demand substantial computational resources [7].
Recently, state-space models have emerged as a solution to mitigate the computational burden of transformers. Their complexity scales subquadratically with the sequence length [8], which allows them to process millions of tokens [9]. However, even in state-space models, very long sequences will impose considerable memory and computational demands.
Bolya et al. [10], have shown that the efficiency of ViTs can be substantially improved by merging tokens throughout the transformer architecture. Specifically, they compute similarity scores between tokens and combine them into single tokens through a convex combination. However, they only explore token merging for ViT architectures.
In this empirical study, we for the first time explore token merging within the time series domain. We introduce a novel local token merging algorithm whose computational complexity varies from linear to quadratic, based on the neighborhood considered for each token merge. This allows token merging to scale to long sequences. The algorithm is illustrated in figure 1. Through comprehensive empirical evaluations, we analyze the impact of token merging on various time series transformer models and state-space models. Our key contributions are as follows:
Token merging in time series We present first studies on token merging in time series analysis, exploring its application beyond transformer architectures to include state-space models. For this purpose, we propose a domain-specific token merging algorithm that combines tokens within a local neighborhood around each token. Adjusting the size of the neighborhood allows the complexity of the algorithm to range from linear to quadratic. We show that this adaptability improves upon classical token merging for long-sequence modeling tasks.
Model acceleration Our findings from both pretrained models and those trained with token merging reveal substantial computational savings with only slight reductions in accuracy across five time series transformer architectures and datasets. We also assess five different model sizes per architecture, noting greater relative accelerations in larger models.
Token merging in foundation models Specifically, token merging enhances the throughput of the time series foundation model Chronos by up to , with a marginal quality drop of in relative MSE. Across four out of five datasets, we identify Pareto optimal points where token merging simultaneously boosts throughput and increases accuracy. For example, on the ETTh1 dataset, token merging achieves a acceleration while the MSE improves by .
Token merging patterns Our analysis of token merging’s effects identifies three distinct patterns: 1) a consistent decline in performance with an increasing number of merged tokens, 2) initial improvements in accuracy with few merged tokens followed by a drop as merging increases, and 3) scenarios where accuracy remains unchanged regardless of the token merging rate. We investigate all patterns and find different explanations for all behaviors.
2 Related work
Time series transformers In recent years, many transformer architectures with inductive biases for time series have been proposed, successfully outperforming classical and other deep-learning-based methods in time series forecasting quality like recurrent neural networks [5]. Most of them focus on reducing complexity by modifying the attention mechanism. LogTrans uses LogSparse attention [5] while Informer focuses only on the most relevant queries using ProbSparse attention [11]. Additionally, many architectures adopt decomposition techniques to model trend and seasonal patterns [12, 13, 14, 15]. Autoformer leverages autocorrelation as a sequence-based similarity measure in the attention mechanism [13]. FEDformer uses the frequency domain to model time series effectively [14]. Non-stationary Transformers further mitigate the effect of the time series distribution changing over time [15]. PatchTST embeds subsequences as tokens to capture local semantic information and reduce complexity [16]. Other works apply hierarchical attention [17, 18] or leverage attention between the time series variates to better model multivariate patterns [19, 20].
Due to their success in the vision and NLP domain, transformer-based foundation models have recently emerged for time series, often used in zero-shot settings. Many works focus on training transformers directly on large and diverse time series datasets, usually with billions of tokens [21, 22, 23, 24]. Inspired by the success of foundation models in NLP, the recently proposed Chronos model converts continuous time series data into a fixed vocabulary and is trained on both real-world and synthetic data [7]. Besides, other research branches focus on fine-tuning vision or NLP models for time series [25] and on applying large language models directly on time series data [26].
State-space models Due to the quadratic scaling of the attention mechanism, transformer architectures suffer from significant computational cost when processing very long input sequences. Recently, state-space models have shown promising results in overcoming the quadratic complexity of transformers with respect to input length. Linear state-space layers solve the sequential processing requirement of RNNs through linear state-space representations [27]. The S4 model reduces memory requirements by conditioning the state-space matrix with a low-rank correction [28]. By using implicit convolutions and a data-aware gating mechanism, Hyena [8] became one of the first state-space model architectures to match transformers on NLP tasks. Later work uses hardware-aware algorithms to improve the performance of state-space models on modern accelerators [29].
Reducing tokens Many works reduce the number of processed tokens to increase the efficiency of transformer architectures in computer vision and NLP, often by pruning [30, 31]. Marin et al. [32] merge tokens in ViT architectures to reduce the loss of information associated with pruning. Bolya et al. [10] enhance the token merging algorithm, which they successfully apply to already trained models. Besides initial work on classification tasks [10], subsequent work applies token merging to diffusion models [33]. Kim et al. [34] combine merging and pruning, while other work investigates optimal merging and pruning rates [35, 36].
3 Token merging
Despite recent advances in efficient transformers, processing long input sequences still induces considerable memory requirements and computational effort. In the following, we first outline the token merging algorithm introduced by Bolya et al. [10] for ViT architectures. We then propose local merging enabling efficient token merging for state-space models and long sequence processing.
(Global) Token merging in computer vision
Let a neural network consist of layers denoted as , where each layer takes the output of the previous layer as input. We assume that the input consists of tokens with dimension . Thereby, the input tokens are generated by a tokenizer out of -dimensional input data . In the computer vision domain, usually takes the form , where are the width, height, and channels of the input image, respectively, and .
To improve the computational efficiency of a given model, Bolya et al. [10] combine the most similar tokens in each layer, reducing the tokens to be processed in layer to . Therefore, they split the set of all tokens into two disjoint subsets in alternation to avoid merging conflicts and allow for a parallelized computation of merging correspondences. Here and contain elements each, denoted as and respectively. The authors compute the cosine similarity between all tokens in both subsets and merge the most similar correspondences by averaging the tokens accordingly. This results in a global token merging algorithm with quadratic complexity. Lastly, the authors use a fixed to enable batch processing without needing to pad individual batch elements to the same shape after token reduction.
(Local) Token merging for time series
In this work, we design token merging mechanisms for time series architectures and demonstrate run-time and even performance improvements over various datasets and models. We assume that the input consists of time stamps with variates.
Previous work on token merging in image processing explored global merging schemes, where every token of each subset and could be merged with each other [10, 33].
However, computing the similarity between both sets of tokens has a complexity of , which is suboptimal for sequential data often consisting of long token sequences [6, 37], and state-space models featuring subquadratic complexity [8, 9].
Therefore, we propose local merging - a superset of token merging - by introducing as a locality constraint where we compute the similarity only on a local subset of tokens . Figure 1 illustrates the proposed merging algorithm. The locality constraint reduces the complexity to . Varying the locality, we achieve linear complexity by considering only neighboring tokens for merging up to quadratic complexity by considering a global merging pool, possibly exploiting more redundancy.
For efficient computation, we refactor into a rectangular tensor. An upper bound for the resulting speed up can be given by . The acceleration of deeper models is expected to increase as more subsequent layers can profit from already merged tokens.
Local merging additionally preserves order and locality as an inductive bias for sequence processing.
Some time series transformers use processing mechanisms that require a minimum number of tokens in the forward pass. To universally enable token merging in these architectures, we further introduce as the minimum number of remaining tokens. When encountering odd numbers of tokens we exclude the most recent token from merging as we expect it to contain the most relevant information.
We derive the complexity of the token merging procedures in section A.1.
Dataset | Layers | Transformer | Autoformer | FEDformer | Informer | Non-stationary | ||||||||||
MSE | Accel. | MSEΔ | MSE | Accel. | MSEΔ | MSE | Accel. | MSEΔ | MSE | Accel. | MSEΔ | MSE | Accel. | MSEΔ | ||
ETTh1 | 2 | |||||||||||||||
4 | ||||||||||||||||
6 | ||||||||||||||||
8 | ||||||||||||||||
10 | ||||||||||||||||
ETTm1 | 2 | |||||||||||||||
4 | ||||||||||||||||
6 | ||||||||||||||||
8 | ||||||||||||||||
10 | ||||||||||||||||
Weather | 2 | |||||||||||||||
4 | ||||||||||||||||
6 | ||||||||||||||||
8 | ||||||||||||||||
10 | ||||||||||||||||
Electricity | 2 | |||||||||||||||
4 | ||||||||||||||||
6 | ||||||||||||||||
8 | ||||||||||||||||
10 | ||||||||||||||||
Traffic | 2 | |||||||||||||||
4 | ||||||||||||||||
6 | ||||||||||||||||
8 | ||||||||||||||||
10 |
4 Experiments
We systematically explore token merging in diverse settings on time series datasets and model architectures in different sizes each. Additionally, we investigate token merging in large foundation models using Chronos in a zero-shot setting [7]. Finally, we demonstrate that token merging can be applied to state-space models for long sequence processing by using a novel local merging algorithm featuring subquadratic complexity.
Datasets We use time series forecasting datasets including ETTh1, ETTm1, Weather, Electricity and Traffic for our transformer experiments. For state-space models, we use the long-range Dummy Mouse Enhancers Ensembl dataset. See section A.2 for more details.
Model architectures For our main experiments, we use architectures, including Autoformer, FEDformer, Informer, Non-stationary Transformer, and the vanilla Transformer [1] as reference. For each model, we evaluate token merging for different model sizes with encoder layers, which we train doing hyperparameter optimization (see section A.2). We use an input length of and prediction horizon samples.
For experiments on Chronos, we use the default input length of and prediction horizon [7]. We compute the median from Chronos probabilistic forecasts and report the MSE.
For our experiments on state-space models, we use HyenaDNA medium, a genomic foundation model [9] based on the Hyena architecture [8]. We use a large input length of nucleotides utilizing Hyenas subquadratic complexity.
Applying token merging We apply token merging between the attention and MLP block in all transformer architectures as in [10]. Additionally, we only merge in the encoder and not in the decoder, as many architectures do not allow for a unified merging scheme. In architectures utilizing additional tensors like attention masks or positional biases, we merge them using the same correspondences. Many transformers exhibit quadratic attention, imposing considerable computational cost. As a result, we do not find the token merging algorithm to introduce a substantially additional overhead. Thus, we choose to profit from a global merging pool for transformer models. In state-space models, we merge tokens after the Hyena operator and choose to not introduce an operation with quadratic complexity into the architecture.
Reproducibility of measurements We report all results on the same Nvidia A6000 GPU. For training, we utilize Nvidia V100 and A100 GPUs (see section A.2). We measure the end-to-end execution time of the models using warm-ups and measurement runs per batch. The standard deviation of the execution time is generally in our experiments. Besides the execution time as practically most relevant quantity, we report FLOPs as a more hardware independent measure using the thop library [38]. We always choose the maximum possible batch size and standardize the results.
5 Results
We first present our main results for token merging on pretrained models and models trained with token merging and subsequently conduct several ablation studies.
5.1 Token merging in pretrained models
We investigate token merging on diverse time series transformer models with different inductive biases. All models are specifically trained on the target dataset and token merging is applied only during inference time, as accelerating already trained models is of high practical relevance. We choose token merging hyperparameters as described in section A.2, selecting the fastest token merging trial on the validation set that is within a increase in MSE compared to the reference without token merging.
If we do not find a trial with token merging satisfying these tight criteria, we report results without token merging, mimicking how token merging might be applied in practice. We perform all selections on the validation set and report all results on the test set.
The vanilla and Non-stationary Transformers have quadratic attention mechanisms, while the remaining architectures feature subquadratic attention complexities of for Autoformer and Informer and for FEDformer.
Regardless, token merging substantially increases the throughput of most models, up to , often with no change in forecasting quality as table 1 shows. In some experiments, token merging even improves the MSE. In line with the formal analysis of potential speed up from token merging conducted in section 3, we generally observe higher accelerations for larger models, as more subsequent layers can profit from already merged tokens.
In some cases, we do not find a model with decent forecasting quality satisfying our criteria. Here, token merging during test only has a larger impact on model accuracy, such as for Autoformer on the Traffic dataset. We address this issue when training with token merging in section 5.2.
5.2 Token merging during training
Here, we apply token merging during training to reduce the model’s sensitivity to the algorithm at inference time. As shown in figure 2, models trained with token merging often outperform those trained without it, even if token merging is not applied during testing. This approach enables us to accelerate models such as Autoformer on the Traffic dataset without sacrificing accuracy, which was previously not feasible when applying token merging only during inference. Additionally, token merging accelerates the training process itself by up to for Autoformer on the Traffic dataset.
Dataset | MSE | Best | Fastest | ||
Accel. | MSEΔ | Accel. | MSEΔ | ||
ETTh1 | |||||
ETTm1 | |||||
Weather | |||||
Electricity | |||||
Traffic |
5.3 Scaling to large models
Foundation models are getting more relevant across domains, including NLP [39], computer vision [40], and time series processing [22].
However, these models have high computational requirements. Thus, we investigate token merging for foundation models on Chronos, a univariate probabilistic model, in zero-shot forecasting setting [7]. We apply token merging during inference only, as training Chronos from scratch is not within the scope of this work.
In all our experiments, we find Pareto optimal points with token merging. For four out of five datasets token merging improves both accuracy and throughput simultaneously (see section A.3). Our results demonstrate that it is often beneficial to choose a larger Chronos model with token merging over a smaller one without as in figure 3.
We report our results in table 2, choosing the best Chronos model without token merging as reference. We illustrate two cases: 1) Selecting the token merging setting that provides the best MSE, 2) selecting the setting with the fastest throughput. For 2), we constrain the MSE of token merging trials to be lower than the second-best model without token merging. In addition, we allow a maximum increase in MSE of compared to the reference. In our experiments we can improve Chronos MSE by up to and speed up inference by .
5.4 Merging patterns
We observe three distinct merging patterns when combining tokens in transformer architectures.
Increasing MSE As the number of merged token increases the MSE increases almost monotonically (see figure 3). This behavior can be explained due to a loss of information when combining multiple tokens and also occurs in the vision domain [10].
Constant MSE For the vanilla Transformer on ETTh1 and for FEDformer on ETTh1, Weather, Electricity, and Traffic, we observe a constant MSE when applying token merging as shown in figure 5. For the Transformer model, we find all tokens to be similar after the first attention block. Thus, token merging does not affect the model performance. Nevertheless, we find that in most cases, these models still provide reasonable forecasts. In our experiments, transformer models trained on larger or more complex datasets containing more variates do not show this behavior. We argue that this might be a limitation of transformers on small time series datasets [41, 42]. Still, token merging successfully improves the throughput while maintaining accuracy for these models.
Decreasing MSE Token merging increases forecasting quality, most prominently in Chronos models as in figure 3. We explain this behavior in section 5.5.
5.5 When does token merging improve model performance
In our experiment, applying token merging sometimes improves MSE. Our hypothesis is that averaging similar tokens smoothes the time series, reducing noise and acting as a low-pass filer. To validate our hypothesis, we low-pass filter the input time series using Gaussian kernels without token merging in figure 4. On ETTh1 and Traffic, both token merging and Gaussian filtering improve the MSE. On the Electricity dataset, token merging and Gaussian filtering do not positively impact the MSE. All of these observations are in line with our hypotheses. Applying token merging together with the Gaussian kernel leads to the best results. Other averaging kernels were significantly worse. We show additional results on ETTm1 and Weather in section A.4.
5.6 Dependencies on input length
Token merging effectively reduces the number of tokens in a transformer layer. Here, we explore if we can achieve similar accelerations while maintaining the same prediction quality by varying the number of input samples . For better comparison, we keep the predicted time series snippet fixed and only vary the input sequence.
Our results demonstrate that varying the input length cannot replace token merging (see also section A.5). In figure 6, we investigate input length dependence for two objectives in more detail: First, we explore the token merging setup that leads to the best MSE and compare the results to the model without merging. Here, token merging yields considerable throughput increases while improving predictive quality at the same time. Secondly, we compare the fastest model with token merging, which shows no quality decreases, to a standard model. We find models with token merging to scale favorable to long sequences compared to models without merging.
5.7 Redundancy of input tokens
Token merging exploits similarities in data. Intuitively, the number of tokens that can be merged without affecting predictive performance should depend on the redundancy of the tokens. We explore factors influencing the redundancy of input tokens, including their number and positional embeddings. In the following, we use Autoformer’s time stamp positional embedding for our ablation.
First, we investigate whether scaling the number of input tokens increases average redundancy on the ETTh1 dataset. As demonstrated in figure 7(a), the same relative number of tokens are merged for a given merging threshold, independent of input length. Therefore, we suggest scaling the number of merged tokens in each layer linearly with the input length.
Positional embeddings add information about the location of a token within a sequence. As a result, two identical tokens without positional embeddings may show considerable differences when positional embeddings are included, potentially preventing merging. However, figure 7(a) shows that this effect on token merging is only marginal.
5.8 Dynamic merging
A fixed merging objective allows for batch processing without needing to pad individual time series to the same length. However, it enforces a fixed among independent batch elements, which might not always be optimal. Determining the number of tokens to be merged dynamically using a similarity threshold might increase quality as no dissimilar tokens are combined. Here, we leverage the single-sample case to explore dynamic merging in optimal conditions. From a practical perspective, this case might be relevant for on-device applications like smartphones or automated driving.
In figure 7(b), we compare token merging utilizing a fixed to dynamic merging varying the cosine similarity threshold. Dynamic merging improves quality slightly in most settings. Therefore, we suggest using a fixed merging schedule for batch applications and dynamic merging just for the single-sample case. There is no equivalent to dynamic merging schedules as they are similarity-based and strongly layer-dependent. We report FLOPs as we observe substantial execution overhead in time measurements.
5.9 Token merging in state-space models
Token merging | Accel. | Accuracy |
No merging | 1.00 | 78.9 % |
Local merging | 74.0 % | |
Global merging | 2.93 | 69.4 % |
State-space models can process very long sequences with millions of tokens due to their subquadratic complexity. Our proposed local merging algorithm is specifically designed to match this subquadratic complexity, enabling effective token merging in state-space models. Additionally, it preserves locality and order as inductive bias for sequence processing.
We explore token merging in HyenaDNA. We use a classification task for the experiment, where the data consists of long genomic sequences with nucleotides each. Our local merging with featuring linear complexity and locality bias outperforms global merging with and quadratic complexity. Table 3 illustrates that local merging achieves substantially larger speed up and better accuracy than global merging. This experiment indicates that architecture and domain-specific biases are important when applying token merging to a task.
Local merging accelerates HyenaDNA with a decrease in accuracy. To the best of our knowledge, this is the first study that investigates merging individual states in state-space models to improve their sequence modeling performance.
6 Conclusion
In this work, we explore token merging in the time series domain for the first time. We conduct an extensive empirical study on transformer architectures and state-space models in diverse settings using various models and datasets. We demonstrate that token merging can successfully accelerate pretrained models and sometimes even improve their prediction quality. We further introduce a domain-specific local merging algorithm with variable complexity and illustrate its effectiveness on the Hyena model. On the long-range Dummy Mouse Enhancers Ensembl dataset, this method outperforms traditional token merging approaches in terms of throughput and accuracy.
Finally, we conduct several ablation studies to investigate when token merging is most effective, including sequence length, positional embedding, and single-sample inference settings.
We hope that token merging will have a positive effect on reducing the resource consumption and environmental impact of time series models.
Limitations
In our work, we restrict the merging algorithm to the encoder part of all transformer architectures. Moreover, we divide all tokens into two sets and restrict merging to occur only between tokens from different sets. Future work can explore more flexible merging schemes for time series-specific architectures. Moreover, we do not conduct ablations on all possible hyperparameters due to the large number of architectures and datasets evaluated in this work. Lastly, we focus primarily on common metrics like MSE and accuracy and do not explore the impact of token merging on other factors, such as robustness to distribution shifts or adversarial attacks [43, 44].
Disclaimer
The results, opinions and conclusions expressed in this publication are not necessarily those of Volkswagen Aktiengesellschaft.
Appendix A Appendix
Supplementary material such as derivations, further details and additional results are listed below.
A.1 Derivations
In the following we derive our theoretical results in section 3.
Complexity of local merging
To compute for local merging we need to compute the main diagonal of and depending on also secondary diagonals which are symmetrical but shorter than the main diagonal for . We derive the complexity of local merging depending on in the following:
Merging speed up bound
We roughly estimate the upper bound of the speed up we can achieve by merging tokens in a -layer transformer model. Therefore, we only consider attention due to its quadratic scaling with . We disregard additional effects reducing speed up such as merging overhead to estimate the upper bound. Further, we assume merging half of the tokens each layer. The attention in the first layer is unaffected by merging as we apply token merging between the attention and MLP.
A.2 Experiments
Here we list additional information concerning our experimental settings and resources.
Datasets We base our experiments on commonly used multivariate time series datasets covering different forecasting applications: ETTh1 and ETTm1 consist of variates measuring the power load and temperature of electric transformers in hourly and quarter-hourly granularity [11].
Weather consists of meteorological quantities such as air temperature and is recorded every minutes in .111https://www.bgc-jena.mpg.de/wetter/
Electricity measures the energy demand of consumers every hour [6].
Traffic consists of sensors in the San Francisco Bay Area measuring the road occupancy hourly [6].
We use the same data splits for training, validation and test as [13] for consistency.
Since the Chronos foundation model operates univariately and requires considerable computational resources, we randomly sample the same time series from the test set for all Chronos evaluations. For the ETTh1 dataset we do not observe relevant differences when comparing the results to the full test set.
To explore token merging in an additional sequence-based domain and on a second task, we use the Dummy Mouse Enhancers Ensembl dataset [37] for classifying genomic data. It contains very long sequences of nucleotides from a mouse.
Hyperparameter optimization For each transformer architecture, model size, and dataset we train models without token merging doing hyperparameter tuning of learning rate and dropout using HEBO [45]. Here we apply token merging during inference-time only. We choose the best model based on its validation MSE. We train models with the found hyperparameters, the minimum possible , and different uniformly spaced until all tokens are merged. We again choose the best model based on the MSE for further evaluation. We do hyperparameter optimization trials of both chosen models, trained with and without token merging, using HEBO to find token merging inference hyperparameters and on the validation set. Finally, we evaluate once on the test set to report our results.
Hyperparameters In table 4 we list the most relevant hyperparameters we used for training the transformer models including the vanilla Transformer, Autoformer, FEDformer, Informer and Non-stationary Transformer. For training and testing HyenaDNA [9] and for testing Chronos [7] we used their default hyperparameters.
Hyperparameter | Value |
Training | |
Seed | 2024 |
Optimizer | Adam [46] |
Learning rate | Search space loguniform[, ] |
Learning rate decay | Exponential, |
Dropout | Search space uniform[0.0, 0.25] |
Batch size | 32 |
Epochs | 100 |
Early stopping patience | 7 |
Loss | MSE |
Model | |
Input length | |
Prediction horizon | |
Token dimension | |
Encoder layers | |
Decoder layers | 1 |
Attention heads | 8 |
MLP hidden dimension | 2048 |
Activation | GELU |
Computational effort We estimate the computational effort for reproducing our experiments in table 5. Please note that we base some of our experiments on model checkpoints acquired in previous experiments.
Experiment | Accelerator | GPU hours |
Token merging in pretrained models | A6000 | 80 |
V100 | 6720 | |
Token merging during training | A6000 | 50 |
V100 | 3840 | |
Scaling to large models | A6000 | 500 |
Token merging improves model performance | A6000 | 30 |
Dependencies on input length | A6000 | 80 |
Redundancy of input tokens | A6000 | 5 |
Dynamic merging | A6000 | 140 |
Token merging in state-space models | A6000 | 40 |
A100 | 6 |
A.3 Scaling to large models
In this section we show complete results on applying token merging to Chronos, a time series foundation model.
A.4 When does token merging improve model performance
We find token merging to have a smoothing effect improving MSE and show our results on all datasets here.
A.5 Dependencies on input length
Here we show an additional evaluation on applying token merging in Chronos models with different input lengths.